# i-LAVA: Insights on Low Latency Voice-2-Voice Architecture for Agents

### v1

@@ I Introduction
With the advent of LLMs and their adoption for business use-cases, several text based Customer Support Channels and other interaction domains have moved to extensive use of AI Agents [3] [10]. The ability of LLMs to natively ingest and output text data as well as multimodal data motivates the widespread use of Agents by ensuring low latency responses powered by the reasoning capabilities [2] [4] . The next frontier to adoption of AI Agents are the channels requiring voice data as input and output. The present state of Multi-Modal LMs lacks the maturity of LLMs in terms of their reasoning and tool-calling capabilities, both essential for an AI Agent [8]. Hence, low latency voice-to-voice models still hold their relevance for industrial use cases viz. customer support voice agents. It is paramount that the Agent is contextually aware to the tone of speech and responds accordingly in a human like voice. We present our work on integrating a Residual Vector Quantization (RVQ) based Text-to-Speech (TTS) model optimized for a low latency pipeline. Our work presents an end-to-end pipeline which takes the above mentioned requirements into account for a low latency voice-to-voice interaction channel.
@@ II Methodology
Our proposed system consists of 2 important blocks: Speech to Text and LLM inference text to speech. The end-to-end pipeline (1) supports live audio transcription followed by LLM augmented response generation and low latency text-to-speech service delivering an interactive real-time experience. The ASR component is configured to start transcribing the live speech chunk by chunk and eventually produce a coherent output which accounts for removal of pauses and gap words. Subsequently the output is passed to the LLM component along with the prior conversation transcriptions and LLM generated text responses for context aware replies. The text output of the LLM is fed to the TTS component which is highly optimized for low latency, delivering the first audio chunk in as low as 640.9 ms in the GPU compute environment. The proposed system has been evaluated over both CPU and GPU environments:
@@ II-1 Automated Speech Recognition
We utilize OpenAI’s whisper v3-large-turbo model [7] for audio transcription. The model is used with the chunked algorithm for faster processing
@@ II-2 LLM
For minimal latency, OpenAI’s gpt-4o-mini was used as the intermediate LLM for a balance between LLM intelligence and latency
@@ II-3 Text to Speech
We utilize the CSM-1B model [1]. The model uses Residual Vector Quantization which implements multiple stages of quantization to encode audio data efficiently. It starts with a base quantizer and then subsequent stages quantize the residual errors from the previous stage, creating a highly accurate audio representation. The model is also capable of processing given text and audio context for improved conversation skills.
@@ II-B End-to-End Pipeline
The end-to-end pipeline described in Figure 1 represents the architecture describing the flow of data and the processes involved. A Voice Activity Detector (Silero-VAD) [9] filters the relevant part of speech and passes it to the ASR component for transcription which is further fed to the LLM for response generation. The context audio, context text and LLM response text are tokenized and concatenated to be fed into the TTS component wherein the tokens are processed through RVQ Iterations. The output of the RVQ process is fed to an Audio Decoder which generates the speech response
@@ III Results
Our experiments on V-2-V architecture demonstrate that all components VAD, ASR, LLM based inference and finally TTS needs to work in tandem to generate real life like streaming conversation. The time profile for the components with the same recorded audio input, with TTS running on 16 RVQ iterations is as follows
@@
ASR (s) 0.5056 2.4502
LLM Generation (s) 2.1502 1.3272
TTS (s) 0.6695 1.6299
@@
III-A TTS Component
First, we present results for which we utilize the same number of Mimi codebooks as the number of RVQ Iterations Table II summarizes the metrics for streaming TTS generation in GPU and CPU compute environments (with 32 RVQ Iterations) with the sentence ”I am an AI, I am designed to assist and provide helpful responses to your queries. I am a machine learning model, trained on a vast amount of text data” being used for benchmarking. GPU compute shows a significant speedup compared to CPU and demonstrates the capability for real-time generation with RTF <1<1 and stable streaming output
Table III and IV summarize the impact of number of RVQ Iterations in both CPU and GPU compute environments. Rows represent various relevant metrics and columns represent the number of RVQ Iterations.
There are common trends observed across both the environments and reflect that decreasing the number of RVQ Iterations improved compute times at the cost of audio quality
@@ III-A TTS Component
The impact of number of RVQ iterations is much more pronounced in the GPU compute environment due to the fact that RVQ iterations involve a lot of sequential computations instructed by standard Python code and PyTorch operations resulting in several handoffs between GPU and CPU computation. This adds to the overhead and the CPU computation bottlenecks the GPU. Next, we present the results of using 32 codebooks with Mimi decoder. Table V presents the quality of the voice outputs in terms of SNR in dB computed using WADA-SNR on GPU compute. While the latency and RTF numbers reamain almost similar to reduced Mimi codebooks supported by the fact that Audio decoding is not one of the bottlenecks of the TTS component, the deciding factor between the possible configurations turns out to be the audio quality.
@@ IV Conclusion
  • LLM response generation time remains a bottleneck for real-time conversation

### v2

@@ I Introduction
With the advent of LLMs and their adoption for business use-cases, several text based Customer Support Channels and other interaction domains have moved to extensive use of AI Agents [3] [12]. Retreival augmented generation based agents powered by Language models have already demonstrated huge promise in knowledge based agents for question answering tasks and [5] [8]. Besides, the ability of LLMs to natively ingest and output text data as well as multimodal data motivates the widespread use of agents by ensuring low latency responses powered by the reasoning capabilities [2] [4]. The next frontier for adoption of AI Agents are the channels requiring voice data as input and output. The present state of Multi-Modal LMs lacks the maturity of LLMs in terms of their reasoning and tool-calling capabilities, both essential for an AI Agent [10]. Hence, low latency voice-to-voice models still hold their relevance for industrial use cases viz. customer support voice agents. It is paramount that the Agent is contextually aware to the tone of speech and responds accordingly in a human like voice. We present our work on integrating a Residual Vector Quantization (RVQ) based Text-to-Speech (TTS) model optimized for a low latency pipeline. Our work presents an end-to-end pipeline which takes the above mentioned requirements into account for a low latency voice-to-voice interaction channel.
@@ II Methodology
Our proposed system consists of 3 important blocks: Speech to Text , LLM inference text to speech and Text to speech (TTS). The end-to-end pipeline (1) supports live audio transcription followed by LLM augmented response generation and low latency text-to-speech service delivering an interactive real-time experience. The ASR component is configured to start transcribing the live speech chunk by chunk and eventually produce a coherent output which accounts for removal of pauses and gap words. Subsequently the output is passed to the LLM component along with the prior conversation transcriptions and LLM generated text responses for context aware replies. The text output of the LLM is fed to the TTS component which is highly optimized for low latency, delivering the first audio chunk in as low as 640.9 ms in the GPU compute environment. The proposed system has been evaluated over both CPU and GPU environments:
@@ II-1 Automated Speech Recognition
We utilized OpenAI’s whisper v3-large-turbo model [9] for audio transcription. The model is used with the chunked algorithm for faster processing.
@@ II-2 LLM
For minimal latency, OpenAI’s gpt-4o-mini was used as the intermediate LLM for a balance between LLM intelligence and latency. We experimented with both streaming the LLM output and one-shot generation. For streaming, we pass the LLM output to TTS every 10 tokens generated.
@@ II-3 Text to Speech
We utilize the CSM-1B model [1]. The model uses Residual Vector Quantization which implements multiple stages of quantization to encode audio data efficiently [7]. It starts with a base quantizer and then subsequent stages quantize the residual errors from the previous stage, creating a highly accurate audio representation. The model is also capable of processing given text and audio context for improved conversation skills .
@@ II-B End-to-End Pipeline
The end-to-end pipeline described in Figure 1 represents the architecture describing the flow of data and the processes involved. A Voice Activity Detector (Silero-VAD) [11] filters the relevant part of speech and passes it to the ASR component for transcription which is further fed to the LLM for response generation. The context audio, context text and LLM response text are tokenized and concatenated to be fed into the TTS component wherein the tokens are processed through RVQ Iterations. The output of the RVQ process is fed to an Audio Decoder which generates the speech response.
@@ III Results
Our experiments on V-2-V architecture demonstrate that all components VAD, ASR, LLM based inference and finally TTS needs to work in tandem to generate real life like streaming conversation.
III-A LLM Streaming
We evaluated our proposed architecture end to end by running all 3 components: ASR, LLM inference and TTS. For the same, we performed time profiling of entire pipleine to measure the latency contributed by the various components with TTS running on 16 RVQ iterations, the results of time profiling are capture in Table 1.
@@ III-A LLM Streaming
LLM Streaming
ASR (s) 0.5056 0.4986 2.4502 7 2.2245
LLM (s) 2.1502 0.1504 1.3272 0.1496
TTS (s) 0.6695 0.6345 1.6299 1.4647
Total (s) 3.3253 1.2835 5.4073 3.8388
Note that the data corresponds to a single instance run on the same input audio. We observe that streaming the LLM response as an input for the TTS component provides significant speedup to the total latency (Total (s) in Table I) experienced between end of speech (as determined by VAD) and playing the first audio chunk generated
III-B TTS Component
@@
First, we present results for which we utilize the same number of Mimi codebooks as the number of RVQ Iterations Table II summarizes the metrics for streaming TTS generation in GPU and CPU compute environments (with 32 RVQ Iterations) with the sentence ”I am an AI, I am designed to assist and provide helpful responses to your queries. I am a machine learning model, trained on a vast amount of text data” with no prior context being used for benchmarking and no LLM streaming. GPU compute shows a significant speedup compared to CPU and demonstrates the capability for real-time generation with RTF <1<1 and stable streaming output.
Table III and IV summarize the impact of number of RVQ Iterations in both CPU and GPU compute environments. Rows represent various relevant metrics and columns represent the number of RVQ Iterations. There are common trends observed across both the environments and reflect that decreasing the number of RVQ Iterations improved compute times at the cost of audio quality.
@@
The impact of number of RVQ iterations is much more pronounced in the GPU compute environment due to the fact that RVQ iterations involve a lot of sequential computations instructed by standard Python code and PyTorch operations resulting in several handoffs between GPU and CPU computation. This adds to the overhead and the CPU computation bottlenecks the GPU. Next, we present the results of using 32 codebooks with Mimi decoder. Table V presents the quality of the voice outputs in terms of SNR in dB computed using WADA-SNR on GPU compute. While the latency and RTF numbers remain almost similar to reduced Mimi codebooks supported by the fact that Audio decoding is not one of the bottlenecks of the TTS component, the deciding factor between the possible configurations turns out to be the audio quality.
@@ IV Conclusion
  • One-shot LLM response generation time remains a bottleneck for real-time conversation
  • Streaming the LLM response substantially decreases the total latency by enabling the TTS component to generate audio chunks parallel to the LLM response generation
  • Time taken by the ASR component may be absorbed into the silence window of the VAD component by enabling chunk based generation to process the audio chunks as they are recorded to essentially result in near zero effective latency contributed by the ASR component
@@
  • First audio chunk latency scales proportionately to size of the audio chunk while maintaining a chunk size agnostic RTF, hence first chunk latency may be optimized by taking streaming input from the LLM component
@@
The ASR generation may also be streamed so that the component may start processing the audio input as soon as it gets a positive from VAD instead of waiting for the entire input audio to be first collected by VAD and then sent to the ASR component. Implementation of ASR streaming would ensure that the ASR processing time is enclosed within the window of silence we set for VAD to detect end of user speech input hence providing 0 latency for the ASR component